Ms. Swanson is a 92-year-old resident at Northern Oaks Nursing Center who is being admitted to the local hospital for dehydration and weight loss of 8 lbs in the past 2 weeks. Her medical history includes advanced dementia, a stroke two years ago that left her unable to feed herself, and high blood pressure. Her anthropometric measurements include: ht 5’2”, wt 91 lbs.

Ms. Swanson’s usual appetite is pretty good. She is on a ground/minced diet due to swallowing difficulties as a result of her stroke. Prior to hospital admission her appetite had declined significantly and she refused to drink most fluids offered to her during the day by nursing and dietary staff.

At the hospital an IV is started to help correct her dehydration (which was confirmed via lab work). Her lab work also indicates her prealbumin is low (15 mg/dL). An inpatient dietitian is assigned to Ms. Swanson to help increase her kcalorie and protein intakes in addition to her overall food intake.

  1. What is Ms. Swanson’s desirable (or healthy) body weight? (show your work)
  2. What is her percent ideal body weight? (%IBW) (show your work)
  3. What does Ms. Swanson’s low prealbumin indicate about her current state of health?
  4. What ideas do you have for increasing Ms. Swanson’s kcalorie and protein intakes?
  5. What ideas do you have for increasing Ms. Swanson’s overall food intake?
Answer #1

●To calculate desirable body weight

=45.5kg+2.3kg for each inch after 5 feet

Here the patient is 5'2" therefore for 2" it is 4.6kg

Now substituting in the formula

=45.5+4.6

=50.1 or 50 kg

The desirable body weight should be 50 kg

●To calculate percent ideal body weight

=Current weight ÷Ideal weight ×100

Here current weight =91lbs (0.454×91=41.3 kg)

Ideal weight =50kg

Substituting in the formula

=41.3÷50×100

=0.826×100

=82.6%

●The low prealbumin levels in an individual indicates that the patient is in malnutrition.

●To increase her calories and protein intake some of the best ways are to make dietary modifications through fluids.Starting a TPN can help to meet all the nutrients lost to be regained and balanced

●The overall idea for Ms .Swanson is to increase the protein intake and calorie intake .